### v1

# EdgeAlpha: Bringing Process Discovery to the Data Sources

Abstract
Process Mining is moving beyond mining traditional event logs and nowadays includes, for example, data sourced from sensors in the Internet of Things (IoT). The volume and velocity of data generated by such sensors makes it increasingly challenging for traditional process discovery algorithms to store and mine such data in traditional event logs. Further, privacy considerations often prevent data collection at a central location in the first place. To address this challenge, this paper introduces EdgeAlpha, a distributed algorithm for process discovery operating directly on sensor nodes and edge devices on a stream of real-time event data. Based on the Alpha Miner, EdgeAlpha tracks each event and its predecessor and successor events directly on the sensor node where the event is sensed and recorded. From this local view, each node in EdgeAlpha derives a partial footprint matrix, which we then merge at a central location, whenever we query the system to compute a process model. EdgeAlpha enables (a) scalable mining, as a node, for each event, only interacts with its predecessors and, when queried, only exchanges aggregates, i.e., partial footprint matrices, with the central location and (b) privacy preserving process mining, as nodes only store their own as well as predecessor and successor events. On the Sepsis Cases event log, for example, a node queries on average 18.7% of all nodes. For the Hospital Log, we can even reduce the overall querying to 3.87% of the nodes.
Keywords:
Distributed Process Mining Event Data Stream In-Network Processing Scalability
1 Introduction
In recent years, process mining has moved beyond traditional event logs and increasingly includes data sourced from IoT sensors, for example, in a smart manufacturing or smart health setting [6, 13, 17]. However, already today, IoT devices generate significantly more data than we can store or transfer to the cloud [8, 15, 24]. This high data velocity is a key challenge for process mining algorithms, which traditionally operate on static event logs, i.e., assume that data is readily available for mining at a central location with vast storage and computing capabilities [3]. Further, privacy considerations and data protection laws often prevent large-scale data collection at a central location in the first place [14, 19, 21]. We argue that the flood of data sourced from spatially distributed IoT devices and concerns for data privacy demand for distributed process discovery directly at the data source.
Figure 1: While the classic Alpha Miner (left) collects all events on a central entity, EdgeAlpha processes events directly at the source (right), and only exchanges partial footprint matrices.
In this paper, we introduce EdgeAlpha, a distributed algorithm for process mining based on the Alpha Miner [4]. Utilizing the constrained compute and storage resources of sensor nodes and edge devices we bring the computation, i.e., process discovery, directly to the devices and eliminate the need for storing event logs at a central location. For each event, that a node in EdgeAlpha senses, it locally stores the event as well as the predecessor and successor of this event and updates a local footprint matrix (FM) accordingly, see Figure 1. Thus, over time, each node constructs a partial FM that represents the partial view of this node on the process. To derive the complete process model, a central entity queries the nodes for their partial FM and combines these to a complete matrix.
To determine the predecessor of an event, the node that records this particular event, has to potentially query all other nodes, which can lead to large communication overheads. In this paper, we show that events of real-world processes statistically only have a few predecessors, which further only rarely change throughout the lifetime of a process. As a result, for each event, a node can very efficiently determine the predecessor, which in turn reduces the need for messaging within the network and thereby increases scalability. Further, we prove that EdgeAlpha guarantees a process model identical to the original Alpha Miner. EdgeAlpha’s approach can be easily adapted to other process mining algorithms that rely on direct successions or FMs, such as the Heuristics Miner [22].
Our dataset analysis reveals that even for datasets with a substantial number of activities, such as the Hospital Log [9] comprising more than 600 distinct activities, the average count of predecessors per activity remains modest, averaging fewer than 7 predecessors. This observation enables EdgeAlpha to reduce the communication overhead for determining predecessor events by up to 96% compared to querying all nodes in the network. Additionally, our findings demonstrate that even datasets with numerous activities, such as the Hospital Log, stabilize in the number of requested nodes after only a relative small number of events. By batching predecessor queries in batches of, for example, 40 events, we show that EdgeAlpha can reduce the average number of queried nodes per event to less than 2.5% of all nodes. In our experiments, we compare EdgeAlpha with a naive baseline that queries all nodes to determine the predecessor as well as a baseline that always queries the correct node immediately.
In summary, the contributions of this paper are as follows.
  • We introduce EdgeAlpha, a distributed Alpha Miner operating directly where the events originate.
  • We analytically and experimentally show the correctness of the algorithm.
  • We show that EdgeAlpha reduces the message count for the predecessor determination by up to 96% compared to the naive approach when querying after each event.
  • We ensure that nodes only share aggregated event data with the central entity.
The remainder of the paper is structured as follows. Section 2 introduces the necessary background. Next, Section 3 presents related work on distributed process discovery. Afterwards, we introduce the design of EdgeAlpha in Section 4. In Section 5, we evaluate EdgeAlpha. We conclude the paper in Section 6.
2 Preliminaries and Background
In this section, we provide the required preliminaries and background. We begin with definitions of direct successions and footprint matrices, and then recapitulate the Alpha Miner Algorithm [4].
We refer to Peterson [20] for details on Petri nets and use standard definitions for activities, event logs, cases, and traces [1, 5]. In the following, let $L$ be an event log over the activity set $T$. We define direct successions and the resulting relationships between two activities following the definitions by Gatta et al. [14].
Definition 1 (Direct Succession)
Activity $b$ directly succeeds $a$ in event log $L$, i.e. $a>_{L}b$, if and only if there exists a trace $\sigma=(t_{1},\dots,t_{n})\in L$ and $j\in\{1,\dots,n-1\}$ such that $t_{j}=a~{}\land~{}t_{j+1}=b$.
Definition 2 (Causality, No Direct Succession)
With the help of Definition 1 we now define causality (1) and no direct succession (2) between two activities.
$a\rightarrow_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$ (1)
$a~{}\#_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a\not>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$ (2)
Next, we recapitulate the fundamentals of the Alpha Miner before later in the paper introducing a distributed variant. First, the algorithm, see Algorithm 1, identifies all activities occurring in the event log $L$ and denotes this set with $T_{L}$. Each activity of this set corresponds to a transition in the final output model computed by the algorithm. Next, the algorithm determines two sets, $T_{I}$ and $T_{O}$, which contain the start activities of the traces in $L$ and the end activities of these traces, respectively. Subsequently, the algorithm generates a set $X_{L}$, which comprises pairs of activity sets where each element of the first set is directly followed by every element of the second set, although not vice versa. Moreover, activities within each set do not directly succeed one another. The Alpha Miner minimizes the set of pairs, followed by the derivation of Petri net places from the minimized pairs in the next step. Finally, the algorithm adds arcs between these places and transitions, forming the resulting Petri net.
Algorithm 1 Alpha Miner
1:$T_{L}~{}=\{t\in T~{}|~{}\exists~{}\sigma\in L:~{}~{}t\in\sigma\}$
2:$T_{I}~{}=\{first(\sigma)\in T~{}|~{}\forall~{}\sigma\in L\}$
3:$T_{O}=\{last(\sigma)\in T~{}|~{}\forall~{}\sigma\in L\}$
[…]
6:$P_{L}=\{p_{(A,B)}~{}|~{}(A,B)\in Y_{L}\}~{}\cup~{}\{i_{L},o_{L}\}$
[…]
8:$\alpha(L)=(P_{L},~{}T_{L},~{}F_{L})$
As part of the calculation of $X_{L}$ in the Alpha Miner, it first computes a so-called footprint matrix (FM). Essentially, the FM serves to store information regarding the direct succession relationships between activities within the event log. For event log $L$, the corresponding footprint matrix $FM$ has dimension $|T_{L}|\times|T_{L}|$, where each activity in $L$ is mapped to a specific row and column in the matrix, i.e., we can enumerate the activities such that each activity number $i$ is element of $\{0,\dots,|T_{L}|-1\}$. This way, $FM_{ij}$ denotes the entry of the footprint matrix that describes the relationship between activities $i$ and $j$.
$FM_{ij}=1\quad$ $\Leftrightarrow\quad i>_{L}j$
$FM_{ij}=0\quad$ $\Leftrightarrow\quad i\not>_{L}j$
For our design, we additionally define partial footprint matrices that contain only one specific column. For $i,j\in\{0,\dots,|T_{L}|-1\}$, we define $FM^{i}$ as the footprint matrix that only contains column $i$ and $FM^{i}_{j}$ as the entry of column $i$ in row $j$.
3 Related Work
Van der Aalst [2] introduces a method for computing partial and overlapping process models in the form of Petri nets from partial event logs and defines a way to merge them to a Petri net of the complete event log. Here, models rather than FMs are created in a distributed manner. Evermann and Assadipour [11] leverage Map-Reduce [7] to parallelize the Alpha Miner, distributing the computational workload across multiple compute nodes. However, their approach still relies on an event log that is initially stored at a central location. Subsequent work [12] extends their approach to the Flexible Heuristics Miner [23] and evaluates the performance of the Map-Reduce based Alpha Miner and the Flexible Heuristics Miner. While these studies focus primarily on optimizing performance and reducing the computation time by distributing the computation load over multiple nodes, our research places dual emphasis on scalability, i.e, processing data directly at the source, and privacy preservation by only storing data aggregates at a central location.
Gatta et al. [14] address privacy concerns in the healthcare domain: Here, each hospital center aggregates their own event data into an FM, such that the others only see the aggregated versions of the data. Just as other works [2, 11, 12], they merge the FMs subsequently. Overall, all these collect event logs at one or a few locations. In EdgeAlpha, in contrast, each node collects a local event log and only shares aggregates, i.e., partial FMs, with the central location. As a result, EdgeAlpha increases scalability and reduces privacy risks.
4 Design
In this section, we first discuss our assumptions before detailing on the design of EdgeAlpha, introducing our strategy to find predecessors and proving correctness.
4.1 Assumptions on Events and Event Logs
In adherence to common practice, we assume events with the attributes case identifier, activity name and timestamp, whereas the case identifier is unique for each case. Further, the original Alpha Miner treats events with identical timestamps within the same case as distinct based on their ordering in the event log. Thus, it produces different outputs for when events with equal timestamps are differently ordered in the event log. As we operate on distributed event logs, we – without loss of generality – assume timestamps within the same case to be distinct to be able to compare the resulting process model of EdgeAlpha to the output of the Alpha Miner and ensure correctness.
4.2 EdgeAlpha Algorithm
First, we define the design setting in which we operate, then illustrate the basic idea before formally introducing EdgeAlpha algorithm. Without loss of generality, and to ease simplicity and comprehensibility, we assume each node to detect exactly one type of activity. Note, that by introducing virtual nodes, we can easily extend the problem setting to allow each node to be responsible for multiple activity types. Furthermore, we assume that all nodes can communicate reliably and without significant delays, i.e., use common communication technologies and protocols such as cellular connectivity, Wi-Fi or Bluetooth.
EdgeAlpha as algorithm has two phases: Phase 1, initiated upon event triggering, and Phase 2, activated upon a process model request. In phase 1, whenever a node records an event, it finds the predecessor of the event by communicating to other nodes in the system and, once it determines the predecessor, it notes the directly follows relation in its local FM. As a result, every node continuously updates its local, partial FM over time. Phase 2 is activated whenever an entity requests a process model. Then all nodes send their partial FM to a central location which then constructs a FM from these partial ones. Overall, phase 2 is activated on demand and triggered by outside events, while phase 1 operates continuously. In the following, we define both phases. For simplicity, we begin by introducing the basic version of Phase 1, which, for each event, sends a message to each node in the system. Later, we introduce a more efficient variant and event batching.
4.2.1 Phase 1 – Event Ordering and Partial Footprint Matrix Construction.
Figure 2: Phase 1 – Event Ordering and Partial Footprint Matrix Construction in EdgeAlpha: Without a central entity, nodes determine the order of events collaboratively using message passing. In this specific setting, event 1 is a start event. Thus, after the node senses the event, it queries for the predecessor event, to which it does not get any responses. Thus, it sets a flag that it detected a start event. For event 2, after sensing the event and querying for the predecessor event, the respective node receives a response including event 1. It increments the value in the corresponding cell of its local FM.
When an event $e$ on node $i$ is triggered, node $i$ executes the following steps, see Figure 2 for an example. First, node $i$ stores event $e$ locally, including its case identifier, the activity name and timestamp. Then, it sends all other nodes a message containing $e$ to determine the predecessor event. When receiving the request from $i$, a node checks its storage for events in the same case that have no successor yet and whose timestamps are smaller than the timestamp of $e$. If such an event exists, the node responds with the potential predecessor event data. If there are multiple responses, node $i$ picks the event with the latest timestamp, that is still smaller than the own one, as predecessor. We denote the predecessor node as node $j$. Node $i$ stores the newly determined predecessor locally, updates its local footprint matrix $FM^{i}$ by increasing the count in the respective cell, and informs $j$. Subsequently, node $j$ saves $e$ as the successor event of its own event.
The case where there are multiple events to choose from, indicates communication delays. If delays result in incorrectly set predecessors or successors, the nodes rectify the stored events once the correct event is identified and update the counts in the successor’s partial FM accordingly. This is why nodes store a partial FM that not only encodes whether a direct succession has occurred, but also count the occurrences.
Each node $k\in\{0,\dots,n-1\}$ only keeps track of its predecessors in its FM. Therefore, each local FM has dimension $n\times 1$, with $n$ the number of activities. The value at index $\ell\in\{0,\dots,n-1\}$ describes how often $k$ directly follows $\ell$ in all of the cases, formally $\ell>_{L}k$. It is necessary to increment $FM^{k}_{\ell}$ by $1$ each time $k$ detects $\ell>_{L}k$ and subtract $1$ if a direct succession gets fixed. If a node does not find a predecessor, it determines that it recorded the start of a trace and logs the event as a start event. A node looks for its predecessor rather than its successor since it can be sure that the preceding event already occurred and, thus, the node can search for it directly.
4.2.2 Phase 2 – Requesting Process Data.
This phase begins when an outside entity, for example a central node, requests a process model. During this phase, we request the local FMs of all nodes and start or end events within any case, see Figure 3. Subsequently, each node responds with its local FM. Before sending, the node changes all entries that are greater than 0 to 1, which aggregates the data even more and pushes it back into the conventional FM structure. To determine whether its respective activity is an end activity, it verifies whether it detected an event lacking a successor. The requesting unit receives all responses, it assembles the footprint matrix $FM^{all}\in\{1,0\}^{n\times n}$ concatenating all columns $FM^{i}$, $i\in\{0,\dots,n-1\}$ of all $n$ nodes, ordered by activity identifier. Further, we merge the start and end activities into the sets $T_{I}$ and $T_{O}$, respectively, following the conventions of the Alpha Miner. Subsequently, the requesting unit proceeds with the remainder of the algorithm, starting from step four, the determination of $X_{L}$, which includes calculating the FM, here $FM^{all}$, see Algorithm 1.
Figure 3: Phase 2 – Requesting Process Data: First, we request the partial FMs and whether the activity was the start or end of any trace from all nodes. Upon receiving the information, we concatenate the matrices, form start and end activity sets, compute the process model.
4.2.3 Most-Frequent-Predecessors (MFPs).
In phase 1 of EdgeAlpha, the event detecting node sends out requests to all other nodes, resulting in $n-1$ requests where $n$ denotes the number of distinct nodes, i.e., activities in the event log. To reduce communication overhead, we search for the predecessor as follows: we query nodes in order of their likeliness to be the predecessor. Thus, when searching for the predecessor of an event, a node first contacts the most likely predecessor and then the second most likely one and so on. For this, each node records how often each predecessor occurs. Given that activities are, in real-world scenarios, preceded by a consistent subset of activities, this approach prioritizes querying the most frequently occurring predecessors. When detecting a start event, the corresponding node has to still query every single node in the network, to be sure that it is the start event. In our evaluation in Section 5 we show, for real-world event logs, that our statistical approach significantly reduces communication cost and thereby ensures the scalability of EdgeAlpha.
4.2.4 Batching.
To further minimize communication overhead in EdgeAlpha, we introduce batching of querying predecessor events. Rather than sending individual requests for each event, nodes accumulate a number of events before dispatching them in a single message. The nodes continue to use MFP Requesting. As soon as the node finds the predecessor of an event, it removes the event from the batch. The node ceases to query additional nodes once the batch is empty.
4.2.5 Correctness.
After introducing the algorithmic design of EdgeAlpha, we next prove its correctness. For this, we show in the following that EdgeAlpha outputs the same process models as the traditional centralized Alpha Miner algorithm.
We begin by introducing notations before defining our claim. Let $FM^{all}$ denote the FM computed from an entire event log $L$ and $n:=|T_{L}|$, where $T_{L}$ denotes all activities in $L$. Further, let $FM^{i}$ for $i\in\{0,\dots,n-1\}$ be the partial FM computed by node $i$, and $L_{i}$ the partial event log that node $i$ has knowledge of. Since $i$ only keeps track of its predecessor events, its footprint matrix $FM^{i}$ contains only one column. In the following correctness claim, we omit start and end activities. The design of EdgeAlpha ensures that a node considers an activity only a start activity or end activity if it corresponds to the start or end event of a full trace. By definition, uniting all these activities into two distinct sets adheres to the Alpha Miner algorithm.
Claim
Merging the partial footprint matrices $FM^{0},\dots,FM^{n-1}$ by concatenating them ordered by activity identifier results in the same FM as the $FM^{all}$, formally $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}~{}=~{}FM^{all}.$
Proof
To prove the claim, we show that all direct successions represented in the partial FMs can be found in $FM^{all}$ and vice versa. Thus, proving the claim of equivalence between the merged partial FMs and the FM with the unrestricted knowledge.
By definition, $FM^{all}\in\{0,1\}^{n\times n}$. As already discussed, for $FM^{i},~{}i\in\{0,\dots,n-1\}$, we assume without loss of generality that $FM^{i}\in\{0,1\}^{n\times 1}$. It directly follows $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}\in\{0,1\}^{n\times n}$.
Let $i>_{L_{j}}j,~{}i,j\in\{0,\dots,n-1\}$ be a direct succession encoded in the merged partial footprint matrix $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}$. Due to the concatenation of the partial matrices, it holds $FM^{j}_{i}=1$. $FM^{j}$ contains all direct successions between its own activity and the activities preceding it. This means, there is a trace in the event log $L$ in which $i$ is followed by $j$. By definition, $FM^{all}$ captures all direct successions in $L$, meaning it must hold $FM^{all}_{ij}=1$.
It remains to show that the opposite direction holds. Let $k>_{L}\ell,~{}k,\ell\in\{0,\dots,n-1\}$ be a direct succession represented in $FM^{all}$, formally $FM^{all}_{k\ell}=1$. Since each node keeps track of all its predecessor activities in its local FM, we know that $FM^{\ell}_{k}=1$. Merging all partial FMs only means concatenating them ordered by activity identifier. Therefore, it follows $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}_{k\ell}=1$. Thus, proving the claim.
$\Box$
4.2.6 Design Discussion.
A major advantage of EdgeAlpha is that event logs are no longer stored in one central place. Instead, in EdgeAlpha, each node stores its local view on the process, i.e, its partial process information, and only sends aggregates (partial FMs) to a central entity. In addition, we use batching, i.e. a message is not sent after every event, but after a certain time or number of events in order to make exact times and case sequences unrecognizable.
Further, the design of EdgeAlpha is directly applicable to process mining algorithms beyond the Alpha Miner. Algorithms that start by building a FM or keep track of direct successions in the event log, can be projected onto our distribution approach. A well-known example of another algorithm is the Heuristics Miner [22]. The vast majority of results that we show in our evaluation in Section 5, such as the efficiency of Phase 1 of EdgeAlpha, is not specific to the Alpha Miner and directly applies, for example, to the Heuristics Miner.
5 Evaluation
In this section, we evaluate EdgeAlpha in terms of performance. We start by introducing our implementation in Section 5.1 and suitable baselines in Section 5.2. In Section 5.3, we introduce on which well-known datasets we conduct the experiments. We analyze these datasets in Section 5.4. In Section 5.5, we examine how MFP Requesting reduces the number of queried nodes for Phase 1 of the algorithm compared to naive querying of all nodes. We continue with an estimation on what FM sizes to expect during an algorithm run in Section 5.6. We finish the section by discussing the overall scalability of EdgeAlpha in Section 5.7
5.1 Implementation
We implement EdgeAlpha in Python and use a network of docker containers to simulate the distributed nature of EdgeAlpha. We also validate that the resulting process model of EdgeAlpha is identical to the one computed by pm4py [^1] https://pm4py.fit.fraunhofer.de using both artificial datasets and common real-world event logs, such as the datasets in Section 5.3.
5.2 Baselines
In our experiments, we compare EdgeAlpha with and without MFP Requesting. Not using MFP Requesting, i.e. querying all nodes (”Query All”), serves as upper bound. On the lower end, a theoretical, oracle-like, approach bounds the number of queried nodes in EdgeAlpha. It requires no requests if the predecessor corresponds to the node itself and otherwise takes exactly one request to identify the predecessor.
EdgeAlpha assumes a setting in which event detection is part of the algorithm, whereas existing algorithms assume the availability of an event log from the beginning. Thus, comparing EdgeAlpha to the traditional Alpha Miner [4] or other distributed Alpha Miners such as the Map-Reduce approach [11] or the Alpha Miner in the healthcare scenario [14] is not viable. The novelty of EdgeAlpha lies in leveraging the event detection phase to pre-process events, a feature absent in other approaches.
5.3 Datasets
We conduct experiments on the datasets Hospital Log [9], Sepsis Cases [18] and an IoT-Enriched Event Log for Process Mining in Smart Factories [^2] We only consider the MainProcess.xes file of the not cleaned version. [17]. These datasets naturally support geographical distribution over multiple nodes. To further substantiate our evaluation, we also include the BPI Challenge 2017 [10] and the Road Traffic Fine Management Process [16]. Table 1 shows general statistics of the datasets, including the number of events, the number of cases, the average case length, and the number of activities. As each node is responsible for a single activity, the dataset, for example, represents IoT networks with size of 11 nodes in case of the Traffic Fine dataset and 624 nodes in case of the Hospital Log.
[…]
Traffic Fine 561,470 150,370 3.73 11
Sepsis 15,214 1,050 14.49 16
Smart Factories 8,607 271 31.76 21
BPIC 2017 1,202,267 31,509 38.16 26
Hospital 150,291 1,143 131.49 624
Table 1: Key data of the datasets showcasing their basic properties such as event counts, the number of cases, the average case length or the number of activities.
Event Log ##\##Start Activities ##\##Self-Loops Avg ##\##pred. (SD) Avg MFP Ratio (SD) Reduction
Traffic Fine 1 4,306 6.36 (2.71) 0.80 (0.17) 64.94%
Sepsis 6 1,034 7.19 (2.67) 0.53 (0.20) 81.30%
Smart Factories 1 5,854 4.62 (3.33) 0.69 (0.07) 92.15%
BPIC 2017 1 448,205 6.85 (5.30) 0.79 (0.18) 93.40%
Hospital 29 26,542 6.78 (12.25) 0.69 (0.28) 96.13%
Table 2: Dataset details that directly influence the number of requests in phase 1 of EdgeAlpha. These include the number of start activities and self-loops as well as details about the predecessors, such as the average number of predecessors or the average ratio of the most frequent predecessor per activity. Additionally, it shows the reduction EdgeAlpha achieves in the average number of queried nodes when using MFP Requesting compared to querying all nodes in the network.
5.4 Dataset Analysis
We first analyze the properties of real-world event logs, see Table 2. For instance, our analysis shows that the Smart Factories event log contains numerous self-loops, which arise due to events having different lifecycle transitions and states in this event log. This characteristic is essential for EdgeAlpha, as self-loops do not require communication. Further, the table shows that activities commonly have only a few predecessors, underlining the potential of MFP Requesting.
5.5 Number of Queried Nodes
To assess EdgeAlpha’s performance, we analyze its two phases separately. The event ordering phase depends highly on the number of potential predecessors a node needs to query. In datasets like the Hospital Log there are up to 623 nodes to query, the network traffic can be considerably large, particularly during concurrent events. MFP Requesting in EdgeAlpha builds on past event data to optimize request in the future. For the first experiments we assume a batch size of one to focus on other characteristics. As we show in the the batch size analysis, the different batch sizes strongly influence the number of queried nodes.
Figure 4: Comparison of average queried nodes with MFP Requesting compared to querying all nodes. EdgeAlpha is able to reduce communication by querying the most frequent predecessors. With knowledge of the start events, EdgeAlpha can save even more communication bandwidth.
Figure 5: Cumulative Distribution Functions illustrating the relationship between the normalized number of requested nodes and the probability of occurrence. This diagram shows the distribution of queried node counts. A steeper curve, such as for the Hospital Log, indicates a higher event proportion necessitating only a small fraction of queried nodes to identify their predecessors, reflecting greater efficiency.
Figure 6: Relationship between batch size and average number of queried nodes per event across all datasets. To be able to compare the datasets with each other, we normalize the average number of queried nodes. Across all datasets, transitioning from a batch size of 1 to 10 results in a more significant reduction in queries than from 40 to 50. With increasing batch size the reduction rate decreases.
Figure 7: Stabilization of queried node count over time in Hospital Log. Since the data is fluctuating, we use moving averages to visualize the stabilization. A reference line (Query All) indicates the maximum number of requested nodes per event. The moving average stabilizes after fewer than 3,000 events. We plot an additional line with a larger window size to show the trend of the stabilization. The plots are limited to the first 6,000 events.
Figure 8: Bar diagram based on the Sepsis Cases illustrating the average number of queried nodes per activity (left y-axis) alongside the number of occurrences of each activity (right y-axis). The x-axis displays the different activities of the event log. To put the averages into relation, we plot the number of requested nodes when querying all nodes as well as the average number of queried nodes over all events. The diagram indicates that the events that query on average the most nodes, occur the least. Start activities are excluded from this conclusion.
5.5.1 Comparison Across Datasets.
Figure 4 illustrates, for each dataset, the average number of nodes queried per event to find the predecessor. Query All denotes the naive way to message every node in the network. This serves as baseline to compare to. With MFP Requesting, nodes require significantly fewer requests to identify their predecessor event. We achieve further message reduction when the process has a designated start event, making predecessor requests for the respective node redundant. Table 2 also summarizes the significant reduction in the number of queried nodes. Notably, this reduction increases with the number of activities, influenced by factors such as an expanding average case length, diminishing significance of start events, varying numbers of self-loops relative to total events, and the presence of just a few highly frequent predecessors.
Figure 6 presents Cumulative Distribution Functions (CDF) comparing all datasets. The Hospital Log shows the steepest increase with over 90% of events requiring queries to no more than 5% of the nodes. Conversely, the Road Traffic Fine Management Process has barely any occurrences where the predecessor can be identified after querying just 5% of the nodes. This deviation stems from the limited number of 11 activities in the Traffic Fine dataset. Hence, within the 5% threshold, only self-loops are considered.
5.5.2 Stabilization Analysis.
We now delve deeper into the specifics of the Hospital Log and the Sepsis Cases log. Figure 7 illustrates how the number of queried nodes stabilizes in the Hospital Log after fewer than 3,000 events. There is still a higher amplitude after 5,000 events. However, there are no long-lasting deviations. This event log serves as a good example of stabilization due to its large activity set. Datasets with more activities take longer to stabilize, as each activity must occur multiple times for the MFPs to be identified.
5.5.3 Activity-wise Analysis.
Figure 8 takes a closer look at the specifics of the Sepsis Cases event log. Depending on the activity under examination, the Sepsis Cases log has varied querying patterns. In contrast to other event logs, the Sepsis Cases do not have a designated start event, but six different ones. However, in about 95% of the cases, a case starts with ”ER Registration”, which explains why the corresponding bar clearly stands out above the others. Although ”Release E” shows an increased query count, it is not a start activity. The increased query count is the result of the limited occurrence (only six times), indicating ongoing stabilization. The situation is similar for the four activities to its left. They occur infrequently, which also explains the higher standard error.
5.5.4 Batch Size Analysis.
Figure 6 illustrates the reduction in the average number of queried nodes per event as the batch size increases. The rate in which the number of queries decreases varies between the datasets. However, for all datasets the rate decreases with a growing batch size. For all datasets, the average number of queried nodes per event already lies under 2.5% nodes in the network with a batch size of 40. For the Traffic Fine datasets, 2.5% are on average 0.25 nodes, whereas for the Hospital Log those are on average 15.6 nodes.
Overall, our results indicate that the reduction in the number of queried nodes depends on the statistical properties of the respective dataset, such as average case lengths, occurrences of self-loops, and the variety of predecessors of each activity. The experiments show that it takes the activities some time to identify their most frequent predecessors and thus to stabilize their average number of queried nodes. While MFP Requesting is suitable for stable processes, adjustments can enhance its suitability for dynamic process changes. One such adaptation involves employing a sliding window approach, where nodes track activity predecessors within a fixed window. This strategy leverages recent data for predictions, amplifying the impact of predecessor changes while concurrently minimizing memory usage.
5.6 Footprint Matrix Sizes
In the second phase of EdgeAlpha, nodes transmit their data to the central entity, generating a fixed number of messages. These messages include the node’s local FM along with two additional Booleans indicating if a node recorded the start or end activity of a trace. Thus, the scalability of these messages depends highly on the size of each partial FM. In EdgeAlpha, each node only keeps track of their predecessors in their FMs, resulting in transmitting FMs across the network of a maximal size of $n\times 1$, where $n$ represents the total number of activities. Even for datasets like the Hospital Log that has an activity count of 624, each node transmits only a list consisting of at most 624 elements. Compared to existing distributed Alpha Miners [11, 14], which often require sending up to quadratic FMs to a central entity, our approach is characterized by a compact message size. Moreover, storing such matrix sizes on IoT nodes is feasible even with constrained resources. For long-running processes, EdgeAlpha can further discard event data related to own, predecessor, and successor events once a node’s partial FM has been collected.
5.7 Evaluation Discussion
EdgeAlpha leverages data locality to enhance its efficiency. Unlike existing distributed Alpha Miners [11, 14] which typically collect the event data completely before processing it and creating process models, EdgeAlpha adopts a proactive strategy by processing data at runtime and at its source. As soon as a process model is requested, EdgeAlpha starts at a much more advanced point in the Alpha Miner algorithm. Consequently, our strategy provides a significant time advantage, especially when handling large datasets with hundreds of activities and thousands of events. Thus, building a process model becomes substantially faster. This approach is particularly beneficial for iterative real-time process model creation or monitoring tasks.
6 Conclusion
In this paper, we introduce EdgeAlpha, a novel distributed approach tailored for process discovery in IoT networks. By decentralizing computations to IoT and edge devices, EdgeAlpha addresses scalability and privacy concerns inherent in centralized approaches. We discuss how the design principles of EdgeAlpha can be applied to other process mining algorithms that construct footprint matrices or record direct successions in the beginning, such as the Heuristics Miner. Moreover, the integration of Most-Frequent-Predecessor Requesting reduces communication overhead, ensuring efficient event handling within the network. The batching of predecessor queries further minimizes communication overhead. Empirical evaluations showcase EdgeAlpha’s immense reduction of messages in real-world scenarios compared to a naive predecessor determination approach. EdgeAlpha achieves a reduction in communication for determining predecessor events by up to 96% compared to querying all network nodes after each event. By batching predecessor queries in groups of 50 events, experiments show an average reduction in the number of queried nodes per event to less than 2.5% of all nodes. We highlight, how EdgeAlpha provides a significant time advantage over algorithms lacking pre-processing capabilities. By avoiding the storage of full event traces and sensitive personal data in central locations, EdgeAlpha ensures data privacy through aggregation that eliminates personal identification.
{credits}
6.0.1 Acknowledgements
This work received funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), grant 496119880.