### v3

# EdgeAlpha: Distributed Process Discovery at

Abstract
Process Mining is moving beyond mining traditional event logs and nowadays includes, for example, data sourced from sensors in the Internet of Things (IoT) in a smart home, factory or hospital setting. The volume and velocity of data generated by such sensors makes it increasingly challenging for traditional process discovery algorithms to store and mine such data in traditional event logs. Further, privacy considerations often prevent data collection at a central location in the first place. To address this challenge, this paper introduces EdgeAlpha, an algorithm for distributed process discovery operating directly on sensor nodes on a stream of real-time event data. Based on the Alpha Miner, EdgeAlpha tracks each event and its predecessor and successor events directly on the sensor node where the event is sensed and recorded. From this local view, each node in EdgeAlpha derives a partial footprint matrix, which we then merge at a central location, whenever we query the system to compute a process model. Our analysis shows that even with over 600 activities, like in the Hospital Log, the average number of predecessors per activity remains below 7. This enables efficient determination of predecessors for each event, reducing the communication overhead by up to 96% compared to querying all nodes and enhancing scalability. Additionally, our analysis demonstrates that the number of queried nodes stabilizes after relatively few events, and batching predecessor queries in groups of 40 reduces the average queried nodes per event to less than 2.5%. EdgeAlpha ensures that the resulting process model is identical to the original Alpha Miner.
Index Terms:
Distributed Process Mining, Event Data Stream, In-Network Processing, Scalability
I Introduction
In recent years, process mining has moved beyond traditional event logs and increasingly includes data sourced from Internet of Things (IoT) sensors, for example, in a smart manufacturing, smart home or smart health setting [1, 2, 3, 4]. However, already today, IoT devices generate significantly more data than we can store or transfer to the cloud [5, 6, 7]. This high data velocity is a key challenge for process mining algorithms, which traditionally operate on static event logs, i.e., assume that data is readily available for mining at a central location with vast storage and computing capabilities [8]. Further, privacy considerations and data protection laws often prevent large-scale data collection at a central location in the first place. For example, patient data in hospitals, personal data of workers, data in smart home setups, or company secrets in cross-organizational processes are commonly not shared for analysis [9, 10, 11]. We argue that the flood of data sourced from spatially distributed IoT devices and concerns for data privacy demand for distributed process discovery directly at the data source.
Figure 1: While the classic Alpha Miner (left) collects all events on a central entity, EdgeAlpha processes them in a distributed fashion directly at the source (right), and only exchanges data aggregates, increasing scalability and privacy.
In this paper, we introduce EdgeAlpha, an algorithm for distributed process mining based on the Alpha Miner [12]. Utilizing the constrained compute and storage resources of sensor nodes, further referred to as nodes, we bring the computation, i.e., process discovery, directly to the devices and eliminate the need for storing event logs at a central location. In contrast to Gatta et al. [9] which enable cross-organizational process mining through the transmission of data aggregates, EdgeAlpha distributes the computation of the footprint matrices (FMs) to the sensor-node level. This approach allows the collection of aggregated data within an organization that would otherwise not be obtainable in raw form due to, for example, too high data velocity or data protection regulations.
For each event that a node in EdgeAlpha detects, it locally stores the event as well as the predecessor and successor of this event and updates a local FM accordingly, see Fig. 1. Thus, over time, each node constructs a partial FM that represents the view of this node on the process. To derive the complete process model, one of possibly multiple (central) entities queries the nodes for their partial FMs and combines these into a complete matrix. EdgeAlpha enables (a) scalable mining, as a node only interacts with its predecessors after each event and, when queried, only exchanges aggregates, i.e., partial FMs, with a central entity and (b) privacy-preserving process mining, as nodes only store their own as well as predecessor and successor events.
To determine an event’s predecessor event, the node that records this particular event potentially has to query all other nodes, which can lead to large communication overheads. In this paper, we show that activities of real-world processes statistically only have a few predecessors activities, which further only rarely change throughout the lifetime of a process. Our dataset analysis reveals that even for datasets with a substantial number of activities, such as the Hospital Log [13] comprising more than 600 distinct activities, the average count of predecessors per activity remains modest, averaging fewer than 7 predecessors. As a result, for each event, a node can very efficiently determine the predecessor reducing the communication overhead by up to 96% compared to querying all nodes in the network and thereby increasing scalability. Additionally, our findings demonstrate that even datasets with numerous activities, such as the Hospital Log, stabilize in the number of requested nodes after only a relative small number of events. Further, by batching predecessor queries in batches of, for example, 40 events, we show that EdgeAlpha can reduce the average number of queried nodes per event to less than 2.5% of all nodes. In our experiments, we compare EdgeAlpha with a naive baseline that queries all nodes to determine the predecessor and a baseline that immediately queries the correct node. Further, we prove that EdgeAlpha guarantees a process model identical to the original Alpha Miner and show that in four out of five datasets an intermediate result FM already complies with the FM of the Alpha Miner by more than 90% after less than 5% of the events of the datasets.
EdgeAlpha’s approach can be easily adapted to other process mining algorithms that rely on direct successions or FMs, such as the Heuristics Miner [14].
In summary, the contributions of this paper are as follows.
  • We introduce EdgeAlpha, a distributed Alpha Miner operating directly where the events originate.
  • We analytically and experimentally show the correctness of the algorithm.
  • We demonstrate that EdgeAlpha’s intermediate results quickly converge to the final result of the Alpha Miner.
  • We show that EdgeAlpha reduces the message count for the predecessor determination by up to 96% compared to the naive approach and by 97.5% when batching queries with a batch size of 40.
  • We ensure that the nodes only share aggregated event data with a central entity.
  • We implement EdgeAlpha and publish its code open-source [^1] Available at: redacted..
The remainder of the paper is structured as follows. Section II introduces the necessary background. Next, Section III presents related work on distributed process discovery. Afterwards, we introduce the design of EdgeAlpha in Section IV. In Section V, we evaluate EdgeAlpha. We conclude the paper in Section VI.
II Preliminaries and Background
In this section, we provide the required preliminaries and background. We provide definitions of direct successions and footprint matrices and recapitulate the steps of the Alpha Miner Algorithm [12] that are relevant to our approach.
We refer to Peterson [15] for details on Petri nets and use standard definitions for activities, event logs, cases, and traces [16, 17]. In the following, let $L$ be an event log over the activity set $T$. We define direct successions and the resulting relationships between two activities following the definitions by Gatta et al. [9].
Definition 1 (Direct Succession)
Activity $b$ directly succeeds $a$ in event log $L$, i.e. $a>_{L}b$, if and only if there exists a trace $\sigma=(t_{0},\dots,t_{n-1})\in L$ and $j\in\{0,\dots,n-2\}$ such that $t_{j}=a~{}\land~{}t_{j+1}=b$.
Definition 2 (Causality, No Direct Succession)
With the help of Definition 1 we now define causality (1) and no direct succession (2) between two activities.
$a\rightarrow_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$ (1)
$a~{}\#_{L}$ $~{}b\!\!\!\!\!\!\!\!\!\!\!\!\!\!$ $\Leftrightarrow\quad~{}a\not>_{L}b~{}~{}\land~{}~{}b\not>_{L}a$ (2)
Next, we recapitulate the initial steps of the Alpha Miner before later in the paper introducing a distributed variant. First, the algorithm identifies all activities occurring in the event log $L$ and denotes this set with $T_{L}$. Then, the algorithm determines two sets, $T_{I}$ and $T_{O}$, which contain the start activities of the traces in $L$ and their end activities, respectively. Subsequently, the algorithm determines all direct successions in $L$ which can be stored in an FM. This FM is then the basis for the further calculations in the algorithm and the resulting Petri net.
A FM stores information regarding the direct succession relationships between activities within the event log. For event log $L$, the corresponding footprint matrix $FM$ has dimension $|T_{L}|\times|T_{L}|$, where each activity in $L$ is mapped to a specific row and column in the matrix, i.e., we can enumerate the activities such that each activity number $i$ is an element of $\{0,\dots,|T_{L}|-1\}$. This way, $FM_{ij}$ denotes the entry of the footprint matrix that describes the relationship between activities $i$ and $j$.
$FM_{ij}=1\quad$ $\Leftrightarrow\quad i>_{L}j$
$FM_{ij}=0\quad$ $\Leftrightarrow\quad i\not>_{L}j$
For our design, we define partial footprint matrices containing only one specific column: For $i,j\in\{0,\dots,|T_{L}|-1\}$, we define $FM^{i}$ as the footprint matrix that only contains column $i$ and $FM^{i}_{j}$ as the entry of column $i$ in row $j$.
III Related Work
Van der Aalst [18] introduces a method for computing partial and overlapping process models in the form of Petri nets from partial event logs and defines a way to merge them to a Petri net of the complete event log. Here, models rather than FMs are created in a distributed manner. Evermann and Assadipour [19] leverage Map-Reduce [20] to parallelize the Alpha Miner, distributing the computational workload across multiple compute nodes. However, their approach still relies on an event log that is initially stored at a central location. Subsequent work [21] extends their approach to the Flexible Heuristics Miner [22] and evaluates the performance of the Map-Reduce-based Alpha Miner and the Flexible Heuristics Miner. While these studies focus primarily on optimizing performance and reducing the computation time by distributing the computation load over multiple nodes, our research emphasizes scalability, i.e., processing data directly at the source and privacy preservation by only storing data aggregates at a central location.
Gatta et al. [9] address privacy concerns in the healthcare domain: In their work, each hospital center aggregates its own event data into an FM, such that the others only see the aggregated versions of the data. Like other works [18, 19, 21], they subsequently merge the FMs. In contrast, EdgeAlpha distributes the FM computations to the sensor nodes that record the events, enabling data aggregation at an earlier stage. This allows organizations to collect and utilize data in aggregated form that would otherwise be restricted in its raw form due to data protection regulations. As a result of processing events directly at their sources, EdgeAlpha increases scalability and reduces privacy risks.
IV Design
In this section, we first discuss our assumptions before detailing on the design of EdgeAlpha, introducing our strategy to find predecessors and proving correctness.
IV-A Assumptions on Events and Event Logs
In adherence to common practice, we assume events with the attributes case ID, activity name, and timestamp, whereas the case ID is unique for each case. Our evaluation shows these assumptions are both realistic and applicable in numerous real-world scenarios. For example, in a hospital each patient is commonly assigned a fixed ID that can serve as a case ID and the events such as blood tests or ultrasound are predefined activities. Further, the original Alpha Miner treats events with identical timestamps within the same case as distinct based on their ordering in the event log. Thus, it produces different outputs for when events with equal timestamps are differently ordered in the event log. As we operate on distributed event logs, we assume timestamps within the same case to be distinct to be able to compare the resulting process model of EdgeAlpha to the output of the Alpha Miner and ensure correctness.
IV-B EdgeAlpha Setting
First, we define the design setting in which we operate, then illustrate the basic idea before formally introducing the EdgeAlpha algorithm. We employ EdgeAlpha in typical IoT application settings: A network of sensor nodes that are, for example, integrated into machines in a smart factory or smart health setting. These sensors can communicate with each other, either directly or via the Internet. Further, we assume that a membership management allows nodes to gain knowledge of other nodes and how to reach them, as common in IoT applications. Together, these allow a node to find the processor and successor of an event. Finally, we assume that their clocks are well synchronized, using, for example, established clock-synchronization algorithms such as NTP [23]. This is essential, as accurate timestamps ensure proper event ordering. Without loss of generality, and to ease simplicity and comprehensibility, we assume each node detects exactly one type of activity. Note that by introducing virtual nodes, we can easily extend the problem setting to allow each node to be responsible for multiple activity types.
IV-C EdgeAlpha Algorithm: Overview
EdgeAlpha consists of two phases: Phase 1 is initiated upon event triggering, and Phase 2 is activated upon a process model request. In Phase 1, nodes communicate to determine predecessors when recording an event and note their directly-follows relations in their local FMs. As a result, every node continuously updates its local, partial FM. Other than Phase 1, Phase 2 is activated on demand whenever an entity requests a process model. Then, all nodes send their partial FMs to a central location, which constructs an FM. We define both phases, starting with the basic version of Phase 1, which sends a message to all nodes in the system for each event, followed by a more efficient variant and event batching.
IV-D Phase 1 – Event Ordering & Partial Footprint Matrices
Figure 2: Phase 1 – Event Ordering and Partial Footprint Matrix Construction in EdgeAlpha: Without a central entity, nodes determine the order of events collaboratively using message passing. In this specific setting, event 1 is a start event. Therefore, after the node detects the event, it queries for the predecessor event, for which it does not get any responses. Thus, the node sets a flag that it detected a start event. For event 2, after sensing the event and querying for the predecessor event, the respective node receives a response, including event 1. It increments the value in the corresponding cell of its local FM.
When an event $e$ on node $i$ is triggered, node $i$ executes the following steps; see Fig. 2 for an example. First, node $i$ stores event $e$ locally, including its case ID, the activity name and timestamp. Then, it sends all other nodes a message containing $e$ to determine the predecessor event. When receiving the request from $i$, a node checks its storage for events in the same case that have no successor yet and whose timestamps are smaller than the timestamp of $e$. If such an event exists, the node responds with the potential predecessor event data. If there are multiple responses, node $i$ picks the event with the latest timestamp that is still smaller than the own one, as predecessor. We denote the node with the predecessor event as node $j$. Node $i$ stores the newly determined predecessor locally, updates its local footprint matrix $FM^{i}$ by increasing the count in the respective cell, and informs $j$. Subsequently, node $j$ saves $e$ as the successor event of its own event.
If a node has to choose its predecessor from several events, it indicates communication delays which may cause incorrectly set predecessors. However, this is automatically rectified during the run of the algorithm as soon as the node that belongs to the incorrectly set predecessor event receives a request from the node of its correct successor. The node sets the correct successor and informs the node corresponding to the incorrect successor. This one starts the predecessor search again. The node rectifies the stored events once the correct event is identified and updates the counts in its partial FM accordingly. For this reason nodes store a partial FM that not only encodes whether a direct succession has occurred but also counts the occurrences. The rectification process works due to a single start event and because the nodes compare the timestamps of set successor events with the timestamps of incoming predecessor-search-requests. As soon as a node has more than one successor in a single trace the incorrectly set one is informed and an total order of events is established.
Each node $k\in\{0,\dots,n-1\}$ only keeps track of its predecessor activities in its FM. Therefore, each local FM has dimension $n\times 1$, with $n$ the number of activities. The value at index $\ell\in\{0,\dots,n-1\}$ describes how often $k$ directly follows $\ell$ in all of the cases, formally $\ell>_{L}k$. It is necessary to increment $FM^{k}_{\ell}$ by $1$ each time $k$ detects $\ell>_{L}k$ and subtract $1$ if a direct succession gets fixed. If a node does not find a predecessor, it determines that it recorded the start of a trace and logs the event as a start event. A node looks for its predecessor rather than its successor since it can be sure that the preceding event already occurred and, thus, the node can search for it immediately.
EdgeAlpha’s design further incorporates a sliding-window-principle for discarding old data in order to respond to the requirements of the resource-restricted sensor nodes. With this principle, as soon as a certain memory size has been reached or a certain time has been exceeded, the node discards the oldest data when adding new data. However, the sliding window approach only gets beneficial with a vast amount of data. The event logs we consider in the evaluation do not fill enough memory to utilize the sliding window.
IV-E Phase 2 – Requesting Process Data
This phase begins when an outside entity, for example, a central node, requests a process model. During this phase, we request the local FMs of all nodes and start or end events within any case, see Fig. 3. Subsequently, each node responds with its local FM. Before sending, the node changes all entries greater than 0 to 1, aggregating the data even more. This converts the local partial FM back into the conventional FM structure. Each node checks if events without a successor were recorded to determine whether its respective activity is an end activity. The requesting entity receives all responses, and it assembles the footprint matrix $FM^{all}\in\{1,0\}^{n\times n}$ concatenating the columns $FM^{i}$, $i\in\{0,\dots,n-1\}$ of all $n$ nodes, ordered by activity ID. Further, we merge the start and end activities into two sets $T_{I}$ and $T_{O}$, following the conventions of the Alpha Miner. Next, the requesting unit proceeds with the remainder of the Alpha Miner algorithm, starting after the calculation of the FM, here $FM^{all}$.
Figure 3: Phase 2 – Requesting Process Data: We request partial FMs and start/end activity flags from all nodes. Upon receiving the data, we concatenate the matrices, form start and end activity sets, and compute the process model.
IV-F Most-Frequent-Predecessors (MFPs)
In Phase 1 of EdgeAlpha, the event detecting node sends out requests to all other nodes, resulting in $n-1$ requests where $n$ denotes the number of distinct nodes, i.e., activities in the event log. To reduce communication overhead, we search for the predecessor by querying nodes in order of their likelihood of being the predecessor. Thus, when searching for an event’s predecessor, a node first contacts the most likely predecessor, then the second most likely one, and so on. For this, during runtime each node continuously records how often each predecessor occurs. As we show later in a dataset analysis, most activities in real-world scenarios are preceded by a typically stable and small subset of activities, this approach prioritizes querying the most frequently occurring predecessors. However, when detecting a start event, the corresponding node still has to query every node in the network to ensure it is the start event. In our evaluation in Section V we show, for real-world event logs, that our statistical approach significantly reduces communication cost and thereby ensures scalability.
IV-G Batching
To further minimize communication overhead in EdgeAlpha, we introduce batching of querying predecessor events. Rather than sending individual requests for each event, nodes accumulate a number of events before dispatching them in a single message. The nodes continue to use MFP Requesting. As soon as the node finds the predecessor of an event, it removes the event from the batch. The node ceases to query additional nodes once the batch is empty.
IV-H Correctness
After introducing the algorithmic design of EdgeAlpha, we next prove its correctness. For this, we show in the following that EdgeAlpha outputs the same process models as the traditional centralized Alpha Miner algorithm.
We begin by introducing notations before defining our claim. Let $FM^{all}$ denote the FM computed from an entire event log $L$ and $n:=|T_{L}|$, where $T_{L}$ denotes all activities in $L$. Further, let $FM^{i}$ for $i\in\{0,\dots,n-1\}$ be the partial FM computed by node $i$, and $L_{i}$ the partial event log that node $i$ has knowledge of. Since $i$ only keeps track of its predecessor events, its footprint matrix $FM^{i}$ contains only one column. In the following correctness claim, we omit start and end activities. The design of EdgeAlpha ensures that a node considers an activity only a start activity or end activity if it corresponds to the start or end event of a full trace. By definition, uniting all these activities into two distinct sets adheres to the Alpha Miner algorithm.
Claim 1
Merging the partial footprint matrices $FM^{0},\dots,FM^{n-1}$ by concatenating them ordered by activity ID results in the same FM as the $FM^{all}$, formally $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}~{}=~{}FM^{all}.$
Proof 1
To prove the claim, we show that all direct successions represented in the partial FMs can be found in $FM^{all}$ and vice versa. Thus proving the claim of equivalence between the merged partial FMs and the FM with unrestricted knowledge.
By definition, $FM^{all}\in\{0,1\}^{n\times n}$. As already discussed, for $FM^{i},~{}i\in\{0,\dots,n-1\}$, we assume without loss of generality that $FM^{i}\in\{0,1\}^{n\times 1}$. It directly follows $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}\in\{0,1\}^{n\times n}$.
Let $i>_{L_{j}}j,~{}i,j\in\{0,\dots,n-1\}$ be a direct succession encoded in the merged partial footprint matrix $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}$. Due to the concatenation of the partial matrices, it holds $FM^{j}_{i}=1$. $FM^{j}$ contains all direct successions between its activity and its preceding activities. This means there is a trace in the event log $L$ in which $i$ is followed by $j$. By definition, $FM^{all}$ captures all direct successions in $L$, meaning it must hold $FM^{all}_{ij}=1$.
It remains to show that the opposite direction holds. Let $k>_{L}\ell,~{}k,\ell\in\{0,\dots,n-1\}$ be a direct succession represented in $FM^{all}$, formally $FM^{all}_{k\ell}=1$. Since each node keeps track of all its predecessor activities in its local FM, we know that $FM^{\ell}_{k}=1$. Merging all partial FMs only means concatenating them ordered by activity ID. Therefore, it follows $\begin{pmatrix}FM^{0}&~{}\dots~{}&FM^{n-1}\end{pmatrix}_{k\ell}=1$. Thus proving the claim. $\Box$
IV-I Design Discussion
A major advantage of EdgeAlpha is that event logs are no longer stored in one central place. Instead each node stores its local view on the process, i.e, its partial process information, and only sends aggregates (partial FMs) to a central entity. In addition, we use batching, i.e. a message is not sent after every event, but after a certain time or number of events in order to make exact times and case sequences unrecognizable.
Further, the design of EdgeAlpha directly applies to process mining algorithms beyond the Alpha Miner. Algorithms that start by building an FM or keeping track of direct successions in the event log can be projected onto our distribution approach. A well-known example of another algorithm is the Heuristics Miner [14]. The vast majority of results that we show in our evaluation in Section V, such as the efficiency of Phase 1 of EdgeAlpha, are not specific to the Alpha Miner and directly apply, for example, to the Heuristics Miner.
V Evaluation
In this section, we evaluate EdgeAlpha in terms of correctness, accuracy, and performance. We present our implementation in Section V-A and suitable baselines in Section V-B. Section V-C introduces on which well-known datasets we conduct the experiments. We analyze these datasets in Section V-D. In Section V-E we discuss correctness experiments and the compliance between the FMs of EdgeAlpha and the traditional Alpha Miner. We examine how MFP Requesting reduces the number of queried nodes for Phase 1 of the algorithm compared to naive querying in Section V-F. Section V-G estimates FM sizes during an algorithm run and Section V-H discusses the overall scalability of EdgeAlpha.
V-A Implementation
We implement EdgeAlpha in Python and use a network of docker containers to simulate the distributed nature of EdgeAlpha.
V-B Baselines
In our experiments, we compare EdgeAlpha with and without MFP Requesting. ”Query All” represents the upper bound where all nodes are queried, while a theoretical oracle-like approach serves as the lower bound. The latter requires no requests if the predecessor corresponds to the node itself, and only one request to identify the predecessor otherwise.
EdgeAlpha’s design assumes event detection as part of the algorithm, whereas existing algorithms assume the availability of an event log from the beginning. Thus, comparing EdgeAlpha to the traditional Alpha Miner [12] or other distributed Alpha Miners such as the Map-Reduce approach [19] or the Alpha Miner in the healthcare scenario [9] is not viable. The novelty of EdgeAlpha lies in utilizing the event detection phase to pre-process events, a feature absent in other approaches.
V-C Datasets
We conduct experiments on the datasets Hospital Log [13], Sepsis Cases [24] and an IoT-Enriched Event Log for Process Mining in Smart Factories [^2] We only consider the MainProcess.xes file of the not cleaned version. [3]. These datasets naturally support geographical distribution over multiple nodes. To further substantiate our evaluation, we also include the BPI Challenge 2017 [25] and the Road Traffic Fine Management Process [26]. In our setting, each activity in a dataset represents a node in an IoT network.
TABLE I: Dataset Characteristics (Standard deviation in parentheses if not stated otherwise).
[…]
Traffic Fine 561,470 11 (1) 150,370 3.73 4,306 6.36 (2.71) 0.80 (0.17) 64.94%
Sepsis 15,214 16 (6) 1,050 14.49 1,034 7.19 (2.67) 0.53 (0.20) 81.30%
Smart Factories 8,607 21 (1) 271 31.76 5,854 4.62 (3.33) 0.69 (0.07) 92.15%
BPIC 2017 1,202,267 26 (1) 31,509 38.16 448,205 6.85 (5.30) 0.79 (0.18) 93.40%
Hospital 150,291 624 (29) 1,143 131.49 26,542 6.78 (12.25) 0.69 (0.28) 96.13%
V-D Dataset Analysis
We analyze the properties of real-world event logs. Next to basic information such as the numbers of events, activities and cases and the average case length, Table I includes data that directly influences the number of requests in Phase 1. These include the number of start activities and self-loops as well as details about the predecessors, such as the average number of predecessors or the average ratio of the most frequent predecessor per activity. Additionally, it shows the reduction EdgeAlpha achieves in the average number of queried nodes when using MFP Requesting compared to querying all nodes in the network.
Even in a dataset like the Hospital Log with over 600 distinct activities, the average number of preceding activities per activity stays below 7, showcasing the potential of MFP Requesting to reduce the average number of queried nodes per event. The dataset analysis reveals that event logs like the Smart Factories dataset contain numerous self-loops, resulting from different lifecycle transitions and states within the recorded events. This characteristic benefits EdgeAlpha, as self-loops do not require communication. For self-loops, too, only in case of communication delays a self-loop may be set wrongfully for which, however, EdgeAlpha provides a rectification procedure.
V-E Experimental Correctness and Fitness Analysis
As a first step, we validate that the resulting process model of EdgeAlpha is identical to the one computed by pm4py [^3] https://pm4py.fit.fraunhofer.de using both artificial datasets and common real-world event logs, such as the datasets in Section V-C.
Next, we examine how the merged FMs obtained by querying the nodes for their partial FMs after an increasing number of events differ from the intermediate FM produced by the traditional Alpha Miner. We compare them by using the footprint-based fitness value implemented in pm4py. Fig. 7 confirms the full compliance of EdgeAlpha’s resulting FM with the Alpha Miner’s FM across all dataset. It also demonstrates that nearly all datasets reach 90% fitness with less than 5% of events. Here, the Smart Factories dataset is an exception, reaching 90% fitness after processing approximately 41% of the events, although it is considerably small, as shown in Table I. The results indicate that querying interim results can be highly informative, potentially eliminating the need to process all events there are typically present in a common event log before requesting the first process model. Moreover, our analysis suggests that reducing the sliding window for discarding old data would have minimal impact on the resulting process model, but is dependent on the current dataset.
V-F Number of Queried Nodes
To assess EdgeAlpha’s performance, we analyze its two phases separately. The event ordering phase depends highly on the number of potential predecessors a node queries, leading to significant network traffic, particularly during concurrent events, in datasets like the Hospital Log, where up to 623 nodes are queried. MFP Requesting in EdgeAlpha builds on past event data to optimize requests in the future. For the first experiments, we assume a batch size of one to focus on other characteristics. As our batch size analysis shows, the different batch sizes strongly influence the number of queried nodes.
Figure 4: Comparison of average queried nodes with MFP Requesting and querying all nodes. Diagram also shows standard error. EdgeAlpha can reduce communication by querying the most frequent predecessors. With knowledge of the start events, EdgeAlpha can save even more communication bandwidth.
Figure 5: Fitness over time. Intermediate result FMs of the data sets quickly converge into the FM of the traditional Alpha Miner. The BPIC 2017 data set, for example, already has a fitness value of over 90% after 200 events compared with the Alpha Miner FM, which is approx. 0.02% of the events in the event log.
Figure 6: Cumulative Distribution Functions illustrating the relationship between the normalized number of requested nodes and the probability of occurrence. A steeper curve indicates a higher event proportion necessitating only a small fraction of queried nodes to identify their predecessors, reflecting greater efficiency.
Figure 7: Relationship between batch size and average number of queried nodes per event. Normalizing the average queried nodes for comparison. Across all datasets, transitioning from a batch size of 1 to 10 results in a more significant reduction in queries than from 40 to 50. With increasing batch size, the reduction rate decreases.
Figure 8: Stabilization of queried node count over time in Hospital Log. Since the data fluctuates, we use moving averages to visualize the stabilization. A reference line (Query All) indicates the maximum number of requested nodes per event. The moving average stabilizes after fewer than 3,000 events. We plot an additional line with a larger window size to show the trend of the stabilization. The plots are limited to the first 6,000 events.
Figure 9: Bar diagram based on the Sepsis Cases illustrating the average number of queried nodes per activity (left y-axis) alongside the number of occurrences of each activity (right y-axis). The x-axis displays the different activities of the event log and highlights the start activities (S). We plot the number of requested nodes when querying all nodes and the average number of queried nodes over all event. The diagram shows that activities with the highest average queries occur the least, excluding start activities.
V-F1 Comparison Across Datasets
Fig. 4 illustrates, for each dataset, the average number of nodes queried per event to find the predecessor. Query All denotes the naive way to message every node in the network. This serves as a baseline to compare to. With MFP Requesting, nodes require significantly fewer requests to identify their predecessor event. We achieve further message reduction when the process has a designated start event, making predecessor requests for the respective node redundant. Table I shows the significant reduction in the number of queried nodes. This reduction increases with the number of activities, influenced by factors such as an expanding average case length, diminishing significance of start events, varying numbers of self-loops relative to total events, and the presence of just a few highly frequent predecessors.
Fig. 7 shows Cumulative Distribution Functions (CDF) for all datasets. The Hospital Log has the steepest increase with over 90% of events needing queries to no more than 5% of the nodes. Conversely, the Traffic Fine dataset has few instances where the predecessor can be identified after querying just 5% of the nodes. This deviation stems from the limited number of 11 activities in the Traffic Fine dataset. Hence, within the 5% threshold, only self-loops are considered.
V-F2 Stabilization Analysis
We now delve deeper into the specifics of the Hospital Log. Fig. 8 illustrates how the number of queried nodes stabilizes in the Hospital Log after fewer than 3,000 events. There is still a higher amplitude after 5,000 events. However, there are no long-lasting deviations. This event log is a good example of stabilization due to its large activity set. Datasets with more activities take longer to stabilize, as each activity must occur multiple times for the MFPs to be identified.
V-F3 Activity-wise Analysis
Fig. 9 takes a closer look at the specifics of the Sepsis Cases log. Depending on the activity, the Sepsis Cases log shows varied querying patterns. In contrast to other event logs, the Sepsis Cases do not have one designated start event but six different ones. However, about 95% of cases start with ”ER Registration”, which explains why the corresponding bar clearly stands out above the others. Although ”Release E” shows an increased query count, it is not a start activity. The increased query count results from the limited occurrence (only six times), indicating ongoing stabilization. The situation is similar for the four activities to its left. They occur infrequently, which also explains the higher standard error.
V-F4 Batch Size Analysis
Fig. 7 illustrates the reduction in the average number of queried nodes per event as the batch size increases. The rate in which the number of queries decreases varies between the datasets. However, for all datasets the rate decreases with a growing batch size. For all datasets, the average number of queried nodes per event already lies under 2.5% of nodes in the network with a batch size of 40. For the Traffic Fine datasets, 2.5% are on average 0.25 nodes, whereas for the Hospital Log, those are on average 15.6 nodes.
Overall, our results indicate that the reduction in the number of queries depends on the statistical properties of the respective dataset, such as average case lengths, occurrences of self-loops, and the variety of predecessors per activity. The experiments show that it takes the nodes some time to identify their most frequent predecessors and thus to stabilize their average number of queried nodes. While MFP Requesting is suitable for stable processes, adjustments can enhance its suitability for dynamic process changes. One such adaptation involves employing a smaller sliding window, where nodes track activity predecessors only within a fixed window. This strategy leverages recent data for predictions, amplifying the impact of predecessor changes while concurrently minimizing memory usage.
V-G Footprint Matrix Sizes
In the second phase of EdgeAlpha, nodes transmit their data to the central entity, generating a fixed number of messages. These messages include the node’s local FM along with two additional Booleans indicating whether a node recorded the start or end activity of a trace. Thus, the scalability of these messages depends highly on the size of each partial FM. In EdgeAlpha, each node only keeps track of their predecessors in its FM, resulting in transmitting FMs across the network of a maximal size of $n\times 1$, where $n$ represents the total number of activities. Even for datasets like the Hospital Log, which has an activity count of 624, each node transmits only a list consisting of at most 624 elements. Compared to existing distributed Alpha Miners [19, 9], which often require sending up to quadratic FMs to a central entity, our approach is characterized by a compact message size. Moreover, storing such matrix sizes on IoT nodes is feasible even with constrained resources. For long-running processes, EdgeAlpha can further discard event data related to own, predecessor, and successor events once a node’s partial FM has been collected.
V-H Evaluation Discussion
EdgeAlpha leverages data locality to enhance its efficiency. Unlike existing distributed Alpha Miners [19, 9], which typically collect the event data completely before processing it and creating process models, EdgeAlpha adopts a proactive strategy by processing data at runtime and at its source. As soon as a process model is requested, EdgeAlpha starts at a much more advanced point in the Alpha Miner algorithm. Consequently, our strategy provides a significant time advantage, especially when handling large datasets with hundreds of activities and thousands of events. Thus, building a process model becomes substantially faster. This approach is particularly beneficial for iterative real-time process model creation or monitoring tasks.
VI Conclusion
In this paper, we introduce EdgeAlpha, a novel distributed approach tailored to process discovery in IoT networks. By distributing computations to IoT and edge devices, EdgeAlpha addresses scalability and privacy concerns inherent in centralized approaches. We discuss how EdgeAlpha’s design principles can be applied to other process mining algorithms that initially construct footprint matrices or record direct successions, such as the Heuristics Miner [14]. Moreover, integrating Most-Frequent-Predecessor Requesting reduces communication overhead, ensuring efficient event handling within the network. The batching of predecessor queries further minimizes communication overhead. Empirical evaluations showcase EdgeAlpha’s immense reduction of messages in real-world scenarios compared to a naive predecessor determination approach. EdgeAlpha achieves a reduction in communication for determining predecessor events by up to 96% compared to querying all network nodes after each event. By batching predecessor queries in groups of 40 events, experiments show an average reduction in the number of queried nodes per event to less than 2.5% of all nodes. We analytically and experimentally demonstrate the correctness of the algorithm and additionally show how the intermediate results of EdgeAlpha quickly converge to the final result of the Alpha Miner. We highlight how EdgeAlpha provides a significant time advantage over algorithms lacking pre-processing capabilities while still producing the same results as the Alpha Miner. By avoiding the storage of full event traces and sensitive personal data in central locations, EdgeAlpha ensures data privacy through aggregation that eliminates personal identification.